OpenStudy (candy13106):

-x-(13+4x)=-3(5-9x)+2

OpenStudy (anonymous):

aren't you getting tired of these?

OpenStudy (candy13106):

yea its the last one

OpenStudy (anonymous):

i went and watched two episodes of judge judy, came back and you are sill at it do you know how to do this yet ? how quick do you need the answer?

OpenStudy (candy13106):

like now and yes i just get confused

OpenStudy (anonymous):

\[-x-(13+4x)=-3(5-9x)+2\] \[-x-12-4x=-15+27x+2\] is step one

OpenStudy (anonymous):

that be the distributive property then \[-5x-12=27x-13\] is combining like terms

OpenStudy (candy13106):

ok

OpenStudy (anonymous):

well except i made a typo and turned the 13 in to a 12 by accident \[-5x-13=27x-13\] is more like it

OpenStudy (anonymous):

add \(5x\) get \[-13=22x-13\] add \(13\) get \[0=22x\] which makes \(x=0\)

OpenStudy (candy13106):

yep thanks so much for everything u r so smart and what grade r u in

OpenStudy (anonymous):

you totally done now? take a break here is a different version of this tune https://www.youtube.com/watch?v=ks2kBSZ__Fg

OpenStudy (anonymous):

lol grade 75

OpenStudy (candy13106):

okay..........................................................................